HIGH RESOLUTION SPECTRA OF CARBON DIOXIDE CLUSTERS IN THE ν_{3} BAND REGION

\underline{A.R.W. MCKELLAR}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, ON K1A 0R6, Canada; MAHIN AFSHARI, M. DEHGHANY, and N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, 2500 University Dr., N.W., Calgary, AB T2N 1N4, Canada.

There is widespread interest in carbon dioxide clusters from a variety of experimental and theoretical perspectives. But in terms of high resolution spectroscopy, the only definitive information concerns (CO_{2})_{2} and (CO_{2})_{3}. The dimer has a planar slipped parallel geometry with C_{2h} symmetry.^{a} Two isomers are known for the trimer: a planar cyclic form with C_{3h} symmetry^{b} and a sort of “barrel-shaped” form with C_{2} symmetry.^{c} Here we analyze two new bands in the CO_{2} ν_{3} region. The first is a dimer combination band near 2382 cm^{−1} whose assignment raises interesting questions about the intermolecular vibrations of (CO_{2})_{2}.^{d} The second band is a trimer band near 2370 cm^{−1} which is very similar to one we observed previously near 2364 cm^{−1}.^{e} We assign it to a combination involving another out-of-plane vibration of the cyclic trimer. In addition to these newly assigned bands, we also discuss a number of clear and (mostly) well-resolved bands which apparently must belong to (CO_{2})_{N} clusters with N in the range 6 ∼ 15. Although they cannot be precisely assigned at this time, these bands offer intriguing future prospects for learning more about the structures and vibrational dynamics of CO_{2} clusters in a challenging and important size range.

aK.W. Jucks, Z.S. Huang, D. Dayton, R.E. Miller, and W.J. Lafferty, J. Chem. Phys. 86, 4341 (1987); M.A. Walsh, T.H.

England, T.R. Dyke, and B.J. Howard, Chem. Phys. Lett. 142, 265 (1987).

bG.T. Fraser, A.S. Pine, W.J. Lafferty and R.E. Miller, J. Chem. Phys. 87, 1502 (1987).

cM.J. Weida and D.J. Nesbitt, J. Chem. Phys. 105, 10210 (1996).

dH. Chen and J.C. Light, J. Chem. Phys. 112, 5070 (2000).

eM. Dehghany, M. Afshari, N. Moazzen-Ahmadi, and A.R.W. McKellar, J. Chem. Phys. 128, 064308 (2008).